Given: X = 100
The St Dev (sigma sign) is = 7
What is the probability that an item will be below 91 P(X < 91)
What is the probability that an item will be above 111 P(X > 111)
The probability that an item will be between 90 and 100 P(90 < X < 100)

Answer by stanbon(75887)

Given: X = 100
The St Dev (sigma sign) is = 7
What is the probability that an item will be below 91 P(X < 91)
Find the z-score of 91:
z(91)=(91-100)/7=-1.2857
Then P(x<91)=P(z<-1,2857)=0.09927...
----------------------
What is the probability that an item will be above 111 P(X > 111)
Do the same thing with 111
z(111)=(111-100)/7=1.5714
P(x>111)=P(z>1.5714)=0.05804...
---------------------
The probability that an item will be between 90 and 100 P(90 < X < 100)
Find the z-score of 90; the z-score of 100 is zero.
You should be able to do this problem now.
Cheers,
Stan H.